# Emergent field theories from neural networks

### v1

@@ Abstract
We establish a duality relation between Hamiltonian systems and neural network-based learning systems. We show that the Hamilton-Jacobi equations for position and momentum variables correspond to the equations governing the activation dynamics of non-trainable variables and the learning dynamics of trainable variables. The duality is then applied to model various field theories using the activation and learning dynamics of neural networks. For Klein-Gordon fields, the corresponding weight tensor is symmetric, while for Dirac fields, the weight tensor must contain an anti-symmetric tensor factor. The dynamical components of the weight and bias tensors correspond, respectively, to the temporal and spatial components of the gauge field.
@@ 2 Hamiltonian system
which are to be compared with the Hamilton-Jacobi equations:
@@
where $\delta_{ij}$ is the Kronecker delta. These equations would be equivalent to the Hamilton-Jacobi equations:
@@
[…]
@@ 3 Klein-Gordon field
[…]
@@
[…]
In the limit of discretized space (i.e., with lattice vectors ${\bf e}_{1}=(1,0,0)$, ${\bf e}_{2}=(0,1,0)$, ${\bf e}_{3}=(0,0,1)$) and time (i.e., with unit time-step), the Hamilton-Jacobi equations take the following form:
@@
but the discrete second derivative is now evaluated between $t-1$ and $t+1$, as opposed to between $t$ and $t+2$ in equation (3.13). Therefore, the duality between learning systems and field theories is not a map, but a relation — a many-to-one relation.
@@ 4.1 Two types of neurons on one-dimensional lattice
$\frac{\partial}{\partial x^{0}}\phi^{1}(x)-\frac{\partial}{\partial x^{1}}\pi^{2}(x)-m\phi^{2}(x)$ $=$ 00 0
$\frac{\partial}{\partial x^{0}}\pi^{1}(x)+\frac{\partial}{\partial x^{1}}\phi^{2}(x)+m\pi^{2}(x)$ $=$ 00 0 (4.18)
$\frac{\partial}{\partial x^{0}}\phi^{2}(x)+\frac{\partial}{\partial x^{1}}\pi^{1}(x)+m\phi^{1}(x)$ $=$ 00 0
$\frac{\partial}{\partial x^{0}}\pi^{2}(x)-\frac{\partial}{\partial x^{1}}\phi^{1}(x)-m\pi^{1}(x)$ $=$ 00 0
@@
$\gamma^{0}\gamma^{1}+\gamma^{1}\gamma^{0}$ $=$ 00 0 (4.23)
@@ 4.2 Four types of neurons on three-dimensional lattice
[…]
@@ 6 Discussion
In this article, we established a duality relation between classical systems and learning systems, introducing a novel framework to understand the connection between physics and learning. Specifically, we demonstrated that the Hamilton-Jacobi equations governing the dynamics of position and momentum variables correspond to equations governing the activation and learning dynamics of non-trainable and trainable variables, respectively. The duality is not merely formal; it can also serve as a useful tool for modeling, for example, field theories through the activation dynamics of the state of neurons and the learning dynamics of biases in a neural network. Furthermore, this approach offers a conceptual bridge between physical and learning systems, providing a fresh perspective on both domains.

### v2

@@ Abstract
We establish a duality relation between Hamiltonian systems and neural network-based learning systems. We show that the Hamilton’s equations for position and momentum variables correspond to the equations governing the activation dynamics of non-trainable variables and the learning dynamics of trainable variables. The duality is then applied to model various field theories using the activation and learning dynamics of neural networks. For Klein-Gordon fields, the corresponding weight tensor is symmetric, while for Dirac fields, the weight tensor must contain an anti-symmetric tensor factor. The dynamical components of the weight and bias tensors correspond, respectively, to the temporal and spatial components of the gauge field.
@@ 2 Hamiltonian system
which are to be compared with the Hamilton’s equations:
@@
where $\delta_{ij}$ is the Kronecker delta. These equations would be equivalent to the Hamilton’s equations:
@@
$H(\phi,\pi)=\frac{1}{2}\pi^{i}(W-I)_{i}^{\phantom{i}j}\pi_{j}+\sum_{i,j}\left(\int d\phi_{i}\left(B^{-T}B^{-1}\right)_{ij}\frac{\partial{\cal V}\left(\phi\right)}{\partial\phi_{j}}\right)+\sum_{i}F_{i}(\phi_{i}-\sum_{j}B_{i}^{\phantom{i}j}\pi_{j})$ (2.16)
@@ 3 Klein-Gordon field
where $\eta_{\mu\nu}$ is the flat metric with signature $\left(+,-,-,-\right)$ and $\mu,\nu\in\{0,1,2,3\}$. The system can be described in terms of Hamilton’s equations:
@@
${\cal H}[\phi,\beta]=\int dx^{D}\left(\frac{1}{2}\beta({x})^{2}+\frac{1}{2}\sum_{\mu=1}^{3}\left(\frac{\partial\phi({x})}{\partial x_{\mu}}\right)^{2}+\frac{1}{2}m^{2}\phi({x})^{2}\right).$ (3.3)
In the limit of discretized space (i.e., with lattice vectors ${\bf e}_{1}=(1,0,0)$, ${\bf e}_{2}=(0,1,0)$, ${\bf e}_{3}=(0,0,1)$) and time (i.e., with unit time-step), the Hamilton’s equations take the following form:
@@
but the discrete second derivative is now evaluated between t1t-1 and $t+1$, as opposed to between $t$ and $t+2$ in equation (3.13). Therefore, the duality between learning systems and field theories is not a map, but a relation — a many-to-one relation.
@@ 4.1 Two types of neurons on one-dimensional lattice
$\frac{\partial}{\partial x^{0}}\phi^{1}(x)-\frac{\partial}{\partial x^{1}}\pi^{2}(x)-m\phi^{2}(x)$ $=$ $0$
$\frac{\partial}{\partial x^{0}}\pi^{1}(x)+\frac{\partial}{\partial x^{1}}\phi^{2}(x)+m\pi^{2}(x)$ $=$ $0$ (4.18)
$\frac{\partial}{\partial x^{0}}\phi^{2}(x)+\frac{\partial}{\partial x^{1}}\pi^{1}(x)+m\phi^{1}(x)$ $=$ $0$
$\frac{\partial}{\partial x^{0}}\pi^{2}(x)-\frac{\partial}{\partial x^{1}}\phi^{1}(x)-m\pi^{1}(x)$ $=$ $0$
@@
$\gamma^{0}\gamma^{1}+\gamma^{1}\gamma^{0}$ $=$ $0$ (4.23)
@@ 4.2 Four types of neurons on three-dimensional lattice
where $\eta_{\mu\nu}$ is the flat metric with signature $\left(+,-,-,-\right)$ and $\mu,\nu\in\{0,1,2,3\}$. Due to (4.25), the gamma matrices (4.31) satisfy the anti-commutation relation
@@ 6 Discussion
In this article, we established a duality relation between classical systems and learning systems, introducing a novel framework to understand the connection between physics and learning. Specifically, we demonstrated that the Hamilton’s equations governing the dynamics of position and momentum variables correspond to equations governing the activation and learning dynamics of non-trainable and trainable variables, respectively. The duality is not merely formal; it can also serve as a useful tool for modeling, for example, field theories through the activation dynamics of the state of neurons and the learning dynamics of biases in a neural network. Furthermore, this approach offers a conceptual bridge between physical and learning systems, providing a fresh perspective on both domains.